This is a word problem. Sally bought 3 chocolate bars and a pack of gum and paid $1.75. Jake bought 2 chocolate bars and 4 packs of gum and paid $2.00.Find the cost a chocolate bar and the cost of a pack of gum. This is what I have done; 3x + 1y = 1.75; 2x+4y = 2.00.
3x + 1y = 1.75; -8x - 16y = -8.00 (multipling by -4)
17y = -6.25; y= 0.37

3x + 1(0.37)=1.75
3x + 0.37 = 1.75
3x = 1.38
x = 0.46
When Checking; 3(0.46) + 1(0.37)=1.75
1.38 + 0.37 = 1.75
1.75 = 1.75 (true)
Te problem is when I plug these numbers into the second equation it does not come out to 2.00 it comes out as 2.40. Please tell me what I am doing wrong.

Answer by ankor@dixie-net.com(12700)

This is a word problem. Sally bought 3 chocolate bars and a pack of gum and paid $1.75. Jake bought 2 chocolate bars and 4 packs of gum and paid $2.00.
Find the cost a chocolate bar and the cost of a pack of gum.
This is what I have done;
:
3x + 1y = 1.75;
:
2x + 4y = 2.00.
:
3x + 1y = 1.75; multipy this by -4 and add it to the 2nd equation:
:
-12x - 4y = -7
+ 2x + 4y = 2
---------------
-10x + 0y = -5
x = -5/-10
x = +.50 cost of the choc bars
:
I am not quite sure what you were doing, why you mult the 2nd eq by -4 is a puzzle to me. Multiplying the 1st equation by -4 would make sense.
:
Anyway find y using the 1st equation:
3x + y = 1.75
y = 1.75 - 3x
y = 1.75 - 3(.50)
Y = .25 cost of the gum.
:
Check our solutions in the 2nd equation:
:
2x + 4y = 2
2(.5) + 4(.25) = 2
:
I think you had the right idea, just got side-tracked somehow.